OpenStudy (anonymous):

hlp plz 14 x 3/4 A. 42 B. 20 C. 8 2/5 D. 28 3/5

OpenStudy (ccswims):

What do you need help on?

OpenStudy (anonymous):

multiplying that equation

OpenStudy (vianne):

Could you organize it more for me? :P Like A, B, C, D and your question at the top?

OpenStudy (anonymous):

What is the inital question asking?

OpenStudy (vianne):

Either I'm doing something wrong or what... I did \[\frac{ 14 }{ 1 } \times \frac{ 3 }{4 } = \frac{ 42 }{ 4} simplified = 20 \frac{ 1 }{ ? }\]

OpenStudy (vianne):

Ugh... Rewriting.. \[Answer: 20\frac{ 1 }{ 2 }\] Is this correct? :P I thought I replaced the question mark :/